If i invested 10,000 between two stocks, after a year returns were 6% and 5.75%, and total return received was 588.75, how much did i invest in each stock?
Answer by richwmiller(17219)

Total amount of money invested: $10000
x+y=10000,
Total yearly interest for the two accounts is: $588.75
0.06*x+0.0575*y=588.75
x=10000-y
Substitute for x
0.06*(10000-y)+0.0575*y=588.75
Multiply out
600-0.06y+0.0575*y=588.75
Combine like terms.
-0.0025*y=-11.25
Isolate y
y=$ 4500.00 at 5.75%
x=10000-y
Calculate x
x=$ 5500.00 at 6%
Check
0.06*5500+0.0575*4500=588.75
330+258.75=588.75
588.75=588.75
If this statement is TRUE and neither x nor y is negative then all is well